OpenStudy (nicole143):

@mathmale Thank you for your help! I am try to find the real or imaginary solutions of a polynomial equation. I worked it down to what I think is right but got stuck in trying to find factors of 400 that equal -40.

OpenStudy (anonymous):

-20*-20=400 -20 + -20 = -40

OpenStudy (nicole143):

Opps! I meant -41, So sorry.

OpenStudy (anonymous):

show me your full quadratic please

OpenStudy (mathmale):

Let me paraphrase that. We want to find two factors of 400 whose sum is -40. ehuman has just done that. Ooops. Nicole, I don't believe there are any nice, simple factors of 400 whose sum is -41.

OpenStudy (nicole143):

Okay, is another way to find what I am looking for? Would you like me to put up what work I did in trying to solve this?

OpenStudy (anonymous):

won't hurt to check that you are correct before moving on.

OpenStudy (mathmale):

I'm with ehuman in suggesting that he and I could possibly tell you more if we could see the quadratic equation that you're trying to factor. Yes, there is a straightforward method for determining the two factors you want. I'd be glad to go over that.

OpenStudy (nicole143):

Okay, thank you and I will put it up in one moment.

OpenStudy (anonymous):

I'll leave it in the capable hands of mathmale so I can play legos with with my son :) I'll check in soon.

OpenStudy (mathmale):

Nice having a son to play LEGO with! Hope you two enjoy your Lego adventures.

OpenStudy (anonymous):

Same one as in my avatar. He's 2 weeks old in the pic, 3 years old now.

OpenStudy (nicole143):

X^4 – 41x^2 = -400 Plus 400 on both sides X^4 – 41x^2 + 400 = 0 Use a to equal x^2 A^2 – 41a + 400 = 0 (a + or - _ ) (a + or - _ ) = 0 The last bit ^ right there is where I stop trying to find factor that times to 400 and add/subtract to -41 Thank you! @ehuman

OpenStudy (mathmale):

Nicole, are you willing to entertain the possibility that the roots (or zeros) of a^2-41a+400 are not "nice," or, in other words, not integers?

OpenStudy (nicole143):

So I would have to use "i" ?

OpenStudy (mathmale):

Not necessarily. I encourage you to use the quadratic formula to find two possible values for your a (remembering that a=x^2). Would you consider doing this?

OpenStudy (nicole143):

Yes, I'll try right now. Just one moment - I am trying to find the page I was using before(in my book) to see if I possibly had the wrong one.

OpenStudy (nicole143):

Okay, I didn't have the wrong one but you may just know a better way(: Thank you again and would you like me to do the quadratic formal from the point a^2 - 41a + 400 = 0?

OpenStudy (mathmale):

I've done that in the meantime. The roots x happen to be real and positive.

OpenStudy (mathmale):

I'm going to give you a choice: should we talk about how we'd find factors of 400 that sum up to -40, OR continue with the present problem, OR tackle a new one? Your call.

OpenStudy (nicole143):

The present one and the 400 to -41 are the same and I would like to stick with figuring them out

OpenStudy (mathmale):

OK. I've used the quadratic equation to find solutions for a^2-41a+400. As before, the solutions I've found are real and positive. This means that you could take either solution and set it equal to x^2 and find values of x that satisfy the equation. However...those x values would be pretty ugly...no integers!

OpenStudy (mathmale):

So: want to try using the quadratic formula yourself to find those roots of a^2-41a+400, or want to borrow mine? your call.

OpenStudy (nicole143):

Is this a way of doing it?

OpenStudy (nicole143):

I didn't want to keep going with it, if it was wrong but if it is right I'll finish it right now.

OpenStudy (mathmale):

Yes, it is. One error there...under the radical, that b^2 is (-41)^2, not -41^2. If you continue with this you'll get the same answer as I. Smart to finish it so that you'll know how to do it on your own next time.

OpenStudy (mathmale):

That "b^2" really means "(b)^2"...whether b itself is + or -, you must square the whole thing, sign and all.

OpenStudy (nicole143):

Okay - I'll finish it.

OpenStudy (anonymous):

they actually end up factoring, I missed it at 1st.

OpenStudy (nicole143):

Once you get to square root 81 do you take its root 9 and leave it has 9/2?

OpenStudy (nicole143):

*as

OpenStudy (mathmale):

ehuman: have you really found two factors of 400 that sum up to -41? Nicole, I'll be right back.

OpenStudy (nicole143):

Okay.

OpenStudy (mathmale):

I'm afraid I owe you both (ehuman and Nicole) an apology. I tried squaring -41 in my head and got 1641, which is just plain wrong. ehuman helped me to focus on this, with the result that I can now say that under the radical sign we should have 1681-1600.

OpenStudy (mathmale):

Therefore, Nicole, our roots are

OpenStudy (anonymous):

-25*-16 = 400 -25+ -16 = -41

OpenStudy (mathmale):

\[x=\frac{ -41+9 }{2 }\]

OpenStudy (mathmale):

and the other root that stems from a (-) sign in front of the 9. Nicole? What are these 2 roots?

OpenStudy (nicole143):

Wow, you tried doing that in your head? I could never. Would it be 41 and not 41- because "b" is negative to start with?

OpenStudy (mathmale):

Nicole, you're absolutely right. My bad. My very bad.

OpenStudy (anonymous):

+/- 5 and +/-4

OpenStudy (nicole143):

Ha, It's okay. Is one of the x's 25?

OpenStudy (anonymous):

remember your substitution

OpenStudy (mathmale):

YES! in answer to y our question, Nicole. YES! in answer to your reminder, ehuman.

OpenStudy (nicole143):

Okay, so how did @ehuman get the 16?

OpenStudy (mathmale):

So, Nicole, as you've probably already figured out for yourself, and which ehuman has done, our a = 25 = x^2. What are the resulting roots?

OpenStudy (mathmale):

I'll let ehuman explain that, since he obviously knows his stuff. But I'm sticking around.

OpenStudy (anonymous):